What is the X and Y of the equation 5x+y=9 10x-7y=-18 by doing elimination?
Answer by faceoff57(108)

5x + y = 9
10x- 7y = -18
7(5x + y = 9)
35x + 7y = 63
10x - 7y = -18
45x = 45
x = 1
5(1) + y = 9
5 + y = 9
y = 4